# Adaptive Exploration for Multi-Reward Multi-Policy Evaluation

### v2

@@ B.2 Sample Complexity Results
The initial part of the proof follows the same technique as in (Al Marjani et al., 2021; Russo & Proutiere, 2023a), and leverages change of measure arguments (Lai & Robbins, 1985; Garivier & Kaufmann, 2016).
Consider a policy-reward pair $\pi\in\Pi,r\in{\cal R}_{\pi}$. For a confusing model $M_{\pi,r}^{\prime}\in{\rm Alt}_{\pi,r}^{\epsilon}(M)$ consider the log-likelihood ratio up to time $t$ of a sequence of observations $(z_{0},z_{1},\dots)$, with $z_{i}=(s_{i},a_{i})$, under the original MDP $M_{r}$ and an alternative model $M_{\pi,r}^{\prime}\in{\rm Alt}_{\pi,r}(M)$:
$L_{t,\pi,r}=\sum_{n=1}^{t}\log\frac{P(s_{n}|z_{n-1})}{P_{\pi,r}^{\prime}(s_{n}|z_{n-1})}.$
Then, as in (Russo & Proutiere, 2023a), one can show that for all $t\in\mathbb{N}$ the following equality holds
$\mathbb{E}_{M}[L_{t,\pi,r}]$ $=\sum_{s,a}\sum_{t=0}^{\infty}\mathbb{P}_{M}[N_{t}(s,a)\geq t]{\rm KL}_{P|P_{\pi,r}^{\prime}}(s,a),$
@@
for some event ${\cal E}$ that is ${\cal F}_{\tau}$-measurable. Then, define the event =${\cal E}=\{\exists\pi\in\Pi,r\in{\cal R}_{\pi}:\|V_{M_{r}}^{\pi}-\hat{V}_{r}\|_{\infty}>\epsilon\}$. By definition of $(\epsilon,\delta)$-PAC algorithm we have $\mathbb{P}_{M}({\cal E})\leq\delta$. If we can show that $\mathbb{P}_{M_{\pi,r}^{\prime}}({\cal E})\geq 1-\delta$, then we can lower bound ${\rm kl}(\mathbb{P}_{M}({\cal E}),\mathbb{P}_{M_{\pi,r}^{\prime}}({\cal E}))\geq{\rm kl}(\delta,1-\delta)$ by the monotonicity properties of the KL-divergence.
To show $\mathbb{P}_{M_{\pi,r}^{\prime}}({\cal E})\geq 1-\delta$, few steps are needed:
@@
$\stackrel{{\scriptstyle(b)}}{{\geq}}\max_{\bar{\pi}\in\Pi,\bar{r}\in{\cal R}_{\bar{\pi}}}\mathbb{P}_{M_{\pi,r}^{\prime}}(\|V_{M_{\bar{r}}}^{\bar{\pi}}-V_{M_{\bar{r}}^{\prime}}^{\bar{\pi}}\|-\|V_{M_{\bar{r}}^{\prime}}^{\bar{\pi}}-\hat{V}_{\bar{r}}\|_{\infty}>\epsilon),$
$\geq\mathbb{P}_{M_{\pi,r}^{\prime}}(\|V_{M_{r}}^{\pi}-V_{M_{r}^{\prime}}^{\pi}\|-\|V_{M_{r}^{\prime}}^{\pi}-\hat{V}_{r}\|_{\infty}>\epsilon),$
@@
where (a) and (d) follow from the Fréchet inequalities; (b) is an application of the triangle inequality; (c) follows from the fact that $M_{\pi,r}^{\prime}$ is confusing for $(\pi,r)$ and the last inequality from the definition of $(\epsilon,\delta)$-PAC algorithm. Henceforth, for a $(\epsilon,\delta)$-PC algorithm one can conclude that $\mathbb{E}_{M}[L_{\tau,\pi,r}]\geq{\rm kl}(\delta,1-\delta)$.
As the inequality above holds for all $\pi\in\Pi,r\in{\cal R}_{\pi}$, we have
$\inf_{\pi\in\Pi,r\in{\cal R}_{\pi},M_{r}^{\prime}\in{\rm Alt}_{\pi,r}^{\epsilon}(M)}$ $\sum_{s,a}\mathbb{E}_{M}[N_{\tau}(s,a)]{\rm KL}_{M|M_{r}^{\prime}}(s,a)$
$\quad\geq{\rm kl}(\delta,1-\delta).$
Lastly, let $\omega(s,a)=\mathbb{E}_{M}[N_{\tau}(s,a)]/\mathbb{E}_{M}[\tau]$. By optimizing the l.h.s. of the previous inequality over all possible allocations $\omega\in\Omega$, we have $\mathbb{E}_{M}[\tau]T_{\epsilon}^{\star}(M)^{-1}\geq{\rm kl}(\delta,1-\delta)$.
Finally, taking the limit as $\delta\to 0$ yields the result as in (Al Marjani et al., 2021). ∎
@@ B.2.2 Optimality of Behavior Policies
An interesting question is to understand in which cases is a policy optimal for exploration. For instance, consider the MDP in Figure 1 and the single target policy $\pi(\cdot|s)=a_{2}\;\forall s$. For that MDP, it is in general sub-optimal to sample according to such policy, since under $\pi$ state $s_{1}$ becomes transient. In the following lemma we prove that a necessary condition for an optimal exploration policy is to guarantee that states with large deviation gap are visited infinitely often.
@@
Since $s_{c}$, by assumption, is transient under $\pi_{e}$, the expected number of returns of the Markov chain to state $s_{c}$ is bounded, i.e., $\sum_{a}\mathbb{E}_{M}^{\pi}[N_{t}(s_{c},a)]<\infty$ for all $t\geq 1$. In other words, we have that $d^{\pi_{e}}(s_{c})=0$.
@@
But $0=d^{\pi_{e}}(s_{c})=\sum_{a}d^{\pi_{e}}(s_{c},a)$. Which, by non-negativity of $d^{\pi_{e}}$, implies that $d^{\pi_{e}}(s_{c},a)=0$ for every $a$. Hence
@@
$\inf_{M^{\prime}\in{\rm Alt}_{\pi,r}^{\epsilon}}\sum_{s,a}d^{\pi_{e}}(s,a){\rm KL}_{P|P^{\prime}}(s,a)=0,$
@@
$(T_{\epsilon}(d^{\pi_{e}};M))^{-1}=0\quad\forall d^{\pi_{e}}.$
@@
We conclude this section with a sufficient condition on the optimality of a general behavior policy $\pi_{\beta}$, when used to evaluate a single policy $\pi$ on a single reward. What the proposition points out is that $\pi_{\beta}$ needs to induce a sampling distribution $d^{\pi_{\beta}}$ that with large sampling rates where $\|\rho_{r}^{\pi}(s)\|_{\infty}$ is sufficiently large. The extension to multi-reward follows naturally.
@@
$\geq\inf_{M^{\prime}\in{\rm Alt}_{\pi,r}^{\epsilon}(M)}d^{\pi_{\beta}}(\hat{s})\pi_{\beta}(\pi(\hat{s})|\hat{s}){\rm KL}_{P|P^{\prime}}(\hat{s},\pi(\hat{s})),$
$=\inf_{M^{\prime}\in{\rm Alt}_{\pi,r}^{\epsilon}(M)}d^{\pi_{\beta}}(\hat{s},\pi(\hat{s})){\rm KL}_{P|P^{\prime}}(\hat{s},\pi(\hat{s})),$
@@
Hence, to that aim let us define
@@
$d(s)$ $=\sum_{s^{\prime},s^{\prime\prime}}d(s^{\prime})P_{\pi}(s^{\prime}|s^{\prime\prime})P_{\pi}(s^{\prime\prime}|s),$
$\geq\sum_{s^{\prime},s^{\prime\prime}}d(s^{\prime})\epsilon^{2}P_{u}(s^{\prime}|s^{\prime\prime})P_{u}(s^{\prime\prime}|s),$
$=\epsilon^{2}\sum_{s^{\prime}}d(s^{\prime})P_{u}^{2}(s^{\prime}|s)\geq\epsilon^{2}\eta_{2}.$
@@
We prove the theorem for single-policy, since the extension to multi-policy follows immediately.
Start by noting that for all $r\in{\cal R}_{\epsilon}$ we have $\{M_{r}^{\prime}:\|V_{M_{r}}^{\pi}-V_{M_{r}^{\prime}}^{\pi}\|_{\infty}>2\epsilon\}\subseteq\cup_{s}\{M_{r}^{\prime}:|V_{M_{r}}^{\pi}(s)-V_{M_{r}^{\prime}}^{\pi}(s)|>2\epsilon\}$. Using this decomposition, we can show that
$T_{\epsilon}(\omega;M)^{-1}\geq\inf_{r\in{\cal R}_{\epsilon}}$ $\min_{s_{0}}\inf_{M_{r}^{\prime}:|V_{M_{r}}^{\pi}({s_{0}})-V_{M_{r}^{\prime}}^{\pi}({s_{0}})|>2\epsilon}\mathbb{E}_{(s^{\prime},a)\sim\omega}[{\rm KL}_{P|P_{r}^{\prime}}(s^{\prime},a)]$
$=\inf_{r\in{\cal R}_{\epsilon}}\min_{s_{0}}\inf_{M_{r}^{\prime}:|V_{M_{r}}^{\pi}({s_{0}})-V_{M_{r}^{\prime}}^{\pi}({s_{0}})|>2\epsilon}\sum_{s^{\prime},a}\omega(s^{\prime},a){\rm KL}_{P|P_{r}^{\prime}}(s^{\prime},a),$
$=\inf_{r\in{\cal R}_{\epsilon}}\min_{s_{0}}\inf_{M_{r}^{\prime}:|V_{M_{r}}^{\pi}({s_{0}})-V_{M_{r}^{\prime}}^{\pi}({s_{0}})|>2\epsilon}\sum_{s}\omega(s,\pi(s)){\rm KL}_{P|P_{r}^{\prime}}({s},\pi({s})),$
@@
$\left|\Delta P(s,\pi(s))^{\top}V_{M_{r}}^{\pi}\right|=\left|\Delta P(s,\pi(s))^{\top}\rho_{r}^{\pi}(s)]\right|,$
Therefore, using Holder’s inequality and Pinsker’s inequality we obtain
$4\epsilon^{2}$ $\leq\max_{s^{\prime}}\frac{2\gamma^{2}}{(1-\gamma)^{2}}\|\Delta P(s^{\prime},\pi(s^{\prime}))\|_{TV}^{2}\|\rho_{r}^{\pi}(s^{\prime})\|_{\infty}^{2},$
$\leq\max_{s^{\prime}}\frac{\gamma^{2}}{(1-\gamma)^{2}}{\rm KL}_{P|P_{r}^{\prime}}(s^{\prime},\pi(s^{\prime}))\|\rho_{r}^{\pi}(s^{\prime})\|_{\infty}^{2},$
$=\max_{s^{\prime}}\frac{\gamma^{2}}{(1-\gamma)^{2}}{\rm KL}_{P|P_{r}^{\prime}}(s^{\prime},\pi(s^{\prime}))\frac{\omega(s^{\prime},\pi(s^{\prime}))}{\omega(s^{\prime},\pi(s^{\prime}))}\|\rho_{r}^{\pi}(s^{\prime})\|_{\infty}^{2},$
$\leq\max_{s^{\prime}}\frac{\gamma^{2}}{(1-\gamma)^{2}}{\rm KL}_{P|P_{r}^{\prime}}(s^{\prime},\pi(s^{\prime}))\omega(s^{\prime},\pi(s^{\prime}))\max_{s}\frac{\|\rho_{r}^{\pi}(s)\|_{\infty}^{2}}{{\omega(s,\pi(s))}}.$
Therefore $\frac{4\epsilon^{2}(1-\gamma)^{2}}{\gamma^{2}}\min_{s}\frac{{\omega(s,\pi(s))}}{\|\rho_{r}^{\pi}(s)\|_{\infty}^{2}}\leq\max_{s^{\prime}}{\rm KL}_{P|P_{r}^{\prime}}(s^{\prime},\pi(s^{\prime}))\omega(s^{\prime},\pi(s^{\prime}))$, and thus
$\sum_{s}\omega(s,\pi(s)){\rm KL}_{P|P_{r}^{\prime}}({s},\pi({s}))\geq\max_{s^{\prime}}{\rm KL}_{P|P_{r}^{\prime}}(s^{\prime},\pi(s^{\prime}))\omega(s^{\prime},\pi(s^{\prime}))\geq\frac{4\epsilon^{2}(1-\gamma)^{2}}{\gamma^{2}}\min_{s}\frac{{\omega(s,\pi(s))}}{\|\rho_{r}^{\pi}(s)\|_{\infty}^{2}}.$
@@
$T_{\epsilon}(\omega;M)^{-1}$ $\geq\inf_{r\in{\cal R}_{\epsilon}}\min_{s}4\epsilon^{2}\frac{(1-\gamma)^{2}\omega(s,\pi(s))}{\gamma^{2}\|\rho_{r}^{\pi}(s)\|_{\infty}^{2}},$

### v3

@@ B.2 Sample Complexity Results
The overall proof strategy is to derive a set of constraints on the number of visits $N_{t}(s,a)=\sum_{n=1}^{t}\mathbf{1}_{((s_{n},a_{n})=(s,a))}$ to a state-action pair $(s,a)$ that any $(\epsilon,\delta)$-PAC strategy must guarantee. Dividing $\mathbb{E}_{M}[N_{t}(s,a)]$ by $\mathbb{E}_{M}[\tau]$ yields a quantity $\omega(s,a)$ that is a distribution over the state-action space, and as $\delta\to 0$, $\omega$ converges to a stationary distribution. In the last part of the proof we discuss the ergodicity of the chain induced by this stationary distribution.
Part 1: set of constraints through a change of measure. The initial part of the proof follows the same technique as in (Al Marjani et al., 2021; Russo & Proutiere, 2023a), and leverages change of measure arguments (Lai & Robbins, 1985; Garivier & Kaufmann, 2016).
Consider a policy-reward pair $\pi\in\Pi,r\in{\cal R}_{\pi}$. For a confusing model $M_{\pi,r}^{\prime}\in{\rm Alt}_{\pi,r}^{\epsilon}(M)$ consider the log-likelihood ratio up to time $t$ of a sequence of observations $(z_{1},z_{2},\dots)$, with $z_{i}=(s_{i},a_{i})$, under the original MDP $M_{r}$ and an alternative model $M_{\pi,r}^{\prime}\in{\rm Alt}_{\pi,r}^{\epsilon}(M)$:
$L_{t,\pi,r}=\sum_{n=1}^{t}\log\frac{P(s_{n+1}|z_{n})}{P_{\pi,r}^{\prime}(s_{n+1}|z_{n})}.$
Then, as in (Al Marjani & Proutiere, 2021), one can show that for all $t\in\mathbb{N}$ the following equality holds
$\mathbb{E}_{M}[L_{t,\pi,r}]$ $=\sum_{s,a}\sum_{u=1}^{\infty}\mathbb{P}_{M}[N_{t}(s,a)\geq u]{\rm KL}_{P|P_{\pi,r}^{\prime}}(s,a),$
@@
[…]
To show $\mathbb{P}_{M_{\pi,r}^{\prime}}({\cal E})\geq 1-\delta$, a few steps are needed:
@@
$\stackrel{{\scriptstyle(b)}}{{\geq}}\max_{\bar{\pi}\in\Pi,\bar{r}\in{\cal R}_{\bar{\pi}}}\mathbb{P}_{M_{\pi,r}^{\prime}}(\|V_{M_{\bar{r}}}^{\bar{\pi}}-V_{M_{\bar{r}}^{\prime}}^{\bar{\pi}}\|_{\infty}-\|V_{M_{\bar{r}}^{\prime}}^{\bar{\pi}}-\hat{V}_{\bar{r}}\|_{\infty}>\epsilon),$
$\geq\mathbb{P}_{M_{\pi,r}^{\prime}}(\|V_{M_{r}}^{\pi}-V_{M_{r}^{\prime}}^{\pi}\|_{\infty}-\|V_{M_{r}^{\prime}}^{\pi}-\hat{V}_{r}\|_{\infty}>\epsilon),$
@@
where (a) and (d) follow from the Fréchet inequalities; (b) is an application of the triangle inequality; (c) follows from the fact that $M_{\pi,r}^{\prime}$ is confusing for $(\pi,r)$ and the last inequality from the definition of $(\epsilon,\delta)$-PAC algorithm. Henceforth, for a $(\epsilon,\delta)$-PAC algorithm one can conclude that $\mathbb{E}_{M}[L_{\tau,\pi,r}]\geq{\rm kl}(\delta,1-\delta)$.
As the inequality above holds for all $\pi\in\Pi,r\in{\cal R}_{\pi}$, we obtain the following set of constraints:
$\inf_{\pi\in\Pi,r\in{\cal R}_{\pi},M_{r}^{\prime}\in{\rm Alt}_{\pi,r}^{\epsilon}(M)}$ $\sum_{s,a}\mathbb{E}_{M}[N_{\tau}(s,a)]{\rm KL}_{P|P_{r}^{\prime}}(s,a)\geq{\rm kl}(\delta,1-\delta).$
Part 2: optimizing the stationary distribution.
Now, let $n(s,a)=\mathbb{E}_{M}[N_{\tau}(s,a)]$ and $\omega(s,a)=\mathbb{E}_{M}[N_{\tau}(s,a)]/\mathbb{E}_{M}[\tau]$. We rewrite the constraints above as follows
$\mathbb{E}_{M}[\tau]\inf_{\pi\in\Pi,r\in{\cal R}_{\pi},M_{r}^{\prime}\in{\rm Alt}_{\pi,r}^{\epsilon}(M)}\sum_{s,a}\omega(s,a){\rm KL}_{P|P_{r}^{\prime}}(s,a)\geq{\rm kl}(\delta,1-\delta).$
Hence, the optimization problem revolves around optimizing $\omega$ in $\inf_{\pi\in\Pi,r\in{\cal R}_{\pi},M_{r}^{\prime}\in{\rm Alt}_{\pi,r}^{\epsilon}(M)}\sum_{s,a}\omega(s,a){\rm KL}_{P|P_{r}^{\prime}}(s,a)$.
By Lemma 1 in (Al Marjani et al., 2021) we also know that an algorithm that navigates the MDP needs to satisfy the constraints
$\left|\sum_{a}n(s,a)-\sum_{s^{\prime},a^{\prime}}P(s|s^{\prime},a^{\prime})n(s^{\prime},a^{\prime})\right|\leq 1\quad\forall s\in{\cal S}.$
After normalizing $n(s,a)$ by $n(s,a)/\log(1/\delta)$, and using that ${\rm kl}(\delta,1-\delta)\underset{\delta\to 0}{\sim}\log(1/\delta)$, as $\delta\to 0$ we find the following lower bound
$\liminf_{\delta\to 0}\frac{\mathbb{E}_{M}[\tau]}{\log(1/\delta)}\geq\frac{1}{H_{\epsilon}^{\star}(M)},$
where $H^{\star}$ formulates the optimization problem over $\omega$ as follows
$H_{\epsilon}^{\star}(M)\coloneqq$ $\sup_{\omega\in\Delta({\cal S}\times{\cal A})}\inf_{\pi\in\Pi,r\in{\cal R}_{\pi},M_{r}^{\prime}\in{\rm Alt}_{\pi,r}^{\epsilon}(M)}\sum_{s,a}\omega(s,a){\rm KL}_{P|P_{r}^{\prime}}(s,a),$
$\hbox{subject to }\forall s\in{\cal S},\;\sum_{a}\omega(s,a)=\sum_{s^{\prime},a^{\prime}}P(s|s^{\prime},a^{\prime})\omega(s^{\prime},a^{\prime}).$
Part 3: irreducibility of the stationary distribution. Let $\Omega_{0}(M)=\{\omega\in\Delta({\cal S}\times{\cal A}):\sum_{a}\omega(s,a)=\sum_{s^{\prime},a^{\prime}}P(s|s^{\prime},a^{\prime})\omega(s^{\prime},a^{\prime})\;\forall s\in{\cal S}\}$. We note that a solution $\omega^{\star}\in\Omega_{0}(M)$ to the above problem that characterizes $H_{\epsilon}^{\star}(M)$ may not necessarily induce an irreducible chain, i.e., under the induced transition $P_{\omega^{\star}}(s^{\prime}|s)=\frac{\sum_{a}P(s^{\prime}|s,a)\omega^{\star}(s,a)}{\sum_{b}\omega^{\star}(s,b)}$ some states may not be accessible ($s^{\prime}$ is accessible from $s$ if there exists $n\in\mathbb{N}_{+}$ such that $\mathbb{P}_{M}(s_{t+n}=s^{\prime}|s_{t}=s,\omega^{\star})>0$ for all $t$ under $\omega^{\star}$ (Puterman, 2014)). Consider an arbitrary starting state $s_{0}$ (by 3.1). If a state $s_{c}$ is not accessible from $s_{0}$ under $\omega^{\star}$, then the number of visits to that state is bounded, i.e., $\sup_{t\geq 1}\sum_{a}\mathbb{E}_{M}[N_{t}(s_{c},a)]<\infty$.
Under 4.3 there exists $\pi\in\Pi,r\in{\cal R}_{\pi}$ such that we can build a confusing model $M_{r}^{\prime}$. In particular, we can build a confusing model $M_{r}^{\prime}$ as in the proof of Proposition 4.1 (see the “sufficient” part of the proof), satisfying
${\rm kl}(1-\delta,\delta)$ $\leq\sum_{s,a}\mathbb{E}_{M}[N_{\tau}(s,a)]{\rm KL}_{P|P_{r}^{\prime}}(s,a),$
$=\mathbb{E}_{M}[N_{\tau}(s_{c},\pi(s_{c}))]{\rm KL}_{P|P_{r}^{\prime}}(s_{c},\pi(s_{c})),$
$=\mathbb{E}_{M}[N_{\tau}(s_{c},\pi(s_{c}))]\Big{[}(1-P(s_{1}|s_{c},\pi(s_{c})))\log\left(\frac{1}{1-\lambda}\right)$
$\qquad\qquad+P(s_{1}|s_{c},\pi(s_{c}))\log\left(\frac{P(s_{1}|s_{c},\pi(s_{c}))}{\lambda+(1-\lambda)P(s_{1}|s_{c},\pi(s_{c}))}\right)\Big{]},$
$<\infty.$
where $s_{1}\in\operatorname*{arg\,max}_{s\in{\cal S}}|\rho_{r}^{\pi}(s_{c},s)|$ and $\lambda$ is a fixed value in $(0,1)$, independent of $\delta$, taken as in the proof of Proposition 4.1 (in that proof we use $\delta$ instead of $\lambda$, not to be confused with the confidence parameter). The boundedness of the right-hand side comes from the boundedness of $\sum_{a}\mathbb{E}_{M}^{\pi}[N_{t}(s_{c},a)]<\infty$ for all $t\geq 1$ and that $\lambda$ is a fixed value in $(0,1)$. However, as $\delta\to 0$ the left-hand side scales as $\log(1/\delta)\to\infty$, which violates the constraint (and the constraint should hold for all $\pi\in\Pi,r\in{\cal R}_{\pi}$). Since this holds for any pair $(s_{0},s_{c})$, and $s_{0}$ is arbitrary, an optimal solution $\omega^{\star}$ must induce an irreducible chain over $\cal S$}.
Remark: in general, note that the set of confusing models puts constraints on pairs $(s,\pi(s))_{s\in{\cal S},\pi\in\Pi}$, but not necessarily on other state-action pairs. Hence, without further assumptions, we do not have guarantees that an optimal solution yields an irreducible chain. This is in contrast to previous work (Al Marjani et al., 2021; Russo & Proutiere, 2023a; Russo & Vannella, 2024), where constraints also affect pairs $s,a\neq\pi(s)$.
Last step: ergodicity. This last step, while not necessary, shows that any optimal irreducible allocation can be well approximated by a solution $\omega$ satisfying $\omega(s,a)>0$ for all $(s,a)$. We show that any $\omega$ satisfying this latter property forms a dense set in $\Omega_{0}(M)$ (and, since such $\omega(s,a)$ is also irreducible over $\cal S$}, it’s also dense in the set $\Omega_{0}(M)\cap\{\omega:P_{\omega}\hbox{ is irreducible over }{\cal S}\}$).
Consider any $\omega^{\star}\in\Omega_{0}(M)$. Let $\lambda\in(0,1)$ and $\omega^{\prime}\in\Omega(M)$. Define then $\omega_{\lambda}=(1-\lambda)\omega^{\star}+\lambda\omega^{\prime}$.
Then, for any convergent sequence $(\lambda_{k})_{k}$ such that $\lim_{k\to\infty}\lambda_{k}=0$, we have that
$\|\omega^{\star}-\omega_{\lambda_{k}}\|=\lambda_{k}\|\omega^{\star}-\omega^{\prime}\|\to 0.$
Since for every $k$ we have that $\omega_{\lambda_{k}}\in\Omega(M)$, and limit points are in $\Omega_{0}(M)$, it follows that $\Omega(M)$ is dense in $\Omega_{0}(M)$.
@@ B.2.2 Optimality of Behavior Policies
An interesting question is to understand in which cases a policy is optimal for exploration. For instance, consider the MDP in Figure 1 and the single target policy $\pi(\cdot|s)=a_{2}\;\forall s$. For that MDP, it is in general sub-optimal to sample according to such policy, since under $\pi$ state $s_{1}$ becomes transient. In the following lemma we prove that a necessary condition for an optimal exploration policy is to guarantee that states with large deviation gap are visited infinitely often.
@@
Since $s_{c}$, by assumption, is transient under $\pi_{e}$, the expected number of returns of the Markov chain to state $s_{c}$ is bounded, i.e., $\sup_{t\geq 1}\sum_{a}\mathbb{E}_{M}^{\pi_{e}}[N_{t}(s_{c},a)]<\infty$. In other words, we have that $d^{\pi_{e}}(s_{c})=0$.
@@
where $d^{\pi_{e}}(s,a)=d^{\pi_{e}}(s)\pi_{e}(a|s)$ (and $d^{\pi_{e}}(s)=\sum_{a}d^{\pi_{e}}(s,a)$). But $0=d^{\pi_{e}}(s_{c})=\sum_{a}d^{\pi_{e}}(s_{c},a)$. Which, by non-negativity of $d^{\pi_{e}}$, implies that $d^{\pi_{e}}(s_{c},a)=0$ for every $a$. Hence
@@
$\inf_{M^{\prime}\in{\rm Alt}_{\pi,r}^{\epsilon}(M)}\sum_{s,a}d^{\pi_{e}}(s,a){\rm KL}_{P|P^{\prime}}(s,a)=0,$
@@
$(T_{\epsilon}(\omega^{\pi_{e}};M))^{-1}=0\quad\hbox{ for any stationary }\omega^{\pi_{e}}(s,a)=d^{\pi_{e}}(s)\omega(s,a)$
@@
We conclude this section with a sufficient condition on the optimality of a general behavior policy $\pi_{\beta}$, when used to evaluate a single policy $\pi$ on a single reward. What the proposition points out is that $\pi_{\beta}$ needs to induce a sampling distribution $d^{\pi_{\beta}}$ with large sampling rates where $\|\rho_{r}^{\pi}(s)\|_{\infty}$ is sufficiently large. The extension to multi-reward follows naturally.
@@
$),\displaystyle\geq\inf_{M^{\prime}\in{\rm Alt}_{\pi,r}^{\epsilon}(M)}d^{\pi_{\beta}}(\hat{s})\pi_{\beta}(\pi(\hat{s})|\hat{s}){\rm KL}_{P|P^{\prime}}(\hat{s},\pi(\hat{s}$)),
$),\displaystyle=\inf_{M^{\prime}\in{\rm Alt}_{\pi,r}^{\epsilon}(M)}d^{\pi_{\beta}}(\hat{s},\pi(\hat{s})){\rm KL}_{P|P^{\prime}}(\hat{s},\pi(\hat{s}$)),
@@
Hence, to that end, define
@@
$d(s)$ $=\sum_{s^{\prime},s^{\prime\prime}}d(s^{\prime})P_{\pi}(s^{\prime\prime}|s^{\prime})P_{\pi}(s|s^{\prime\prime}),$
$\geq\sum_{s^{\prime},s^{\prime\prime}}d(s^{\prime})\epsilon^{2}P_{u}(s^{\prime\prime}|s^{\prime})P_{u}(s|s^{\prime\prime}),$
$=\epsilon^{2}\sum_{s^{\prime}}d(s^{\prime})P_{u}^{2}(s|s^{\prime})\geq\epsilon^{2}\eta_{2}.$
@@
We prove the theorem considering the following general characteristic time
$T_{\epsilon}(\omega;M)^{-1}\coloneqq\inf_{\pi\in\Pi,r\in{\cal R}_{\pi}^{\epsilon},M_{r}^{\prime}\in{\rm Alt}_{\pi,r}^{\epsilon}(M)}\mathbb{E}_{\omega}[{\rm KL}_{P|P_{r}^{\prime}}(s,a)],$
and we also consider the single-policy case, since the extension to multi-policy follows immediately.
Start by noting that for all $r\in{\cal R}_{\pi}^{\epsilon}$ we have $\{M_{r}^{\prime}:\|V_{M_{r}}^{\pi}-V_{M_{r}^{\prime}}^{\pi}\|_{\infty}>2\epsilon\}\subseteq\cup_{s}\{M_{r}^{\prime}:|V_{M_{r}}^{\pi}(s)-V_{M_{r}^{\prime}}^{\pi}(s)|>2\epsilon\}$. Using this decomposition, we can show that
$T_{\epsilon}(\omega;M)^{-1}\geq\inf_{r\in{\cal R}_{\pi}^{\epsilon}}$ $\min_{s_{0}}\inf_{M_{r}^{\prime}:|V_{M_{r}}^{\pi}({s_{0}})-V_{M_{r}^{\prime}}^{\pi}({s_{0}})|>2\epsilon}\mathbb{E}_{(s^{\prime},a)\sim\omega}[{\rm KL}_{P|P_{r}^{\prime}}(s^{\prime},a)]$
$=\inf_{r\in{\cal R}_{\pi}^{\epsilon}}\min_{s_{0}}\inf_{M_{r}^{\prime}:|V_{M_{r}}^{\pi}({s_{0}})-V_{M_{r}^{\prime}}^{\pi}({s_{0}})|>2\epsilon}\sum_{s^{\prime},a}\omega(s^{\prime},a){\rm KL}_{P|P_{r}^{\prime}}(s^{\prime},a),$
$=\inf_{r\in{\cal R}_{\pi}^{\epsilon}}\min_{s_{0}}\inf_{M_{r}^{\prime}:|V_{M_{r}}^{\pi}({s_{0}})-V_{M_{r}^{\prime}}^{\pi}({s_{0}})|>2\epsilon}\sum_{s}\omega(s,\pi(s)){\rm KL}_{P|P_{r}^{\prime}}({s},\pi({s})),$
@@
$\left|\Delta P(s,\pi(s))^{\top}V_{M_{r}}^{\pi}\right|=\left|\Delta P(s,\pi(s))^{\top}\rho_{r}^{\pi}(s)\right|,$
Therefore, using Holder’s inequality, that $\|p\|_{1}\leq 2\|p\|_{TV}$ for a distribution $p$, and Pinsker’s inequality, we obtain
$4\epsilon^{2}$ $\leq\max_{s^{\prime}}\frac{4\gamma^{2}}{(1-\gamma)^{2}}\|\Delta P(s^{\prime},\pi(s^{\prime}))\|_{TV}^{2}\|\rho_{r}^{\pi}(s^{\prime})\|_{\infty}^{2},$
$\leq\max_{s^{\prime}}\frac{2\gamma^{2}}{(1-\gamma)^{2}}{\rm KL}_{P|P_{r}^{\prime}}(s^{\prime},\pi(s^{\prime}))\|\rho_{r}^{\pi}(s^{\prime})\|_{\infty}^{2},$
$=\max_{s^{\prime}}\frac{2\gamma^{2}}{(1-\gamma)^{2}}{\rm KL}_{P|P_{r}^{\prime}}(s^{\prime},\pi(s^{\prime}))\frac{\omega(s^{\prime},\pi(s^{\prime}))}{\omega(s^{\prime},\pi(s^{\prime}))}\|\rho_{r}^{\pi}(s^{\prime})\|_{\infty}^{2},$
$\leq\max_{s^{\prime}}\frac{2\gamma^{2}}{(1-\gamma)^{2}}{\rm KL}_{P|P_{r}^{\prime}}(s^{\prime},\pi(s^{\prime}))\omega(s^{\prime},\pi(s^{\prime}))\max_{s}\frac{\|\rho_{r}^{\pi}(s)\|_{\infty}^{2}}{{\omega(s,\pi(s))}}.$
Therefore $\frac{2\epsilon^{2}(1-\gamma)^{2}}{\gamma^{2}}\min_{s}\frac{{\omega(s,\pi(s))}}{\|\rho_{r}^{\pi}(s)\|_{\infty}^{2}}\leq\max_{s^{\prime}}{\rm KL}_{P|P_{r}^{\prime}}(s^{\prime},\pi(s^{\prime}))\omega(s^{\prime},\pi(s^{\prime}))$, and thus
$\sum_{s}\omega(s,\pi(s)){\rm KL}_{P|P_{r}^{\prime}}({s},\pi({s}))\geq\max_{s^{\prime}}{\rm KL}_{P|P_{r}^{\prime}}(s^{\prime},\pi(s^{\prime}))\omega(s^{\prime},\pi(s^{\prime}))\geq\frac{2\epsilon^{2}(1-\gamma)^{2}}{\gamma^{2}}\min_{s}\frac{{\omega(s,\pi(s))}}{\|\rho_{r}^{\pi}(s)\|_{\infty}^{2}}.$
@@
$T_{\epsilon}(\omega;M)^{-1}$ $\geq\inf_{r\in{\cal R}_{\pi}^{\epsilon}}\min_{s}2\epsilon^{2}\frac{(1-\gamma)^{2}\omega(s,\pi(s))}{\gamma^{2}\|\rho_{r}^{\pi}(s)\|_{\infty}^{2}},$